OpenStudy (anonymous):

: i need plato help ! What is (5x)(14x3)?

OpenStudy (anonymous):

Ready to work this together?

OpenStudy (anonymous):

yess :) im ready. its for a semester test . so i need help plss ! i would appreciate iht(:

OpenStudy (anonymous):

Sure thing. Just want to double check: 5x times 14x to the 3rd, right?

OpenStudy (anonymous):

Is that the correct interpretation of the question?

OpenStudy (anonymous):

don't bother, some idiot already just gave the answer, so she's moved on..

OpenStudy (anonymous):

Haha oh well. Unfortunate, as these folks will often be the ones who suffer on the tests. :/

OpenStudy (anonymous):

except that I think a lot of them do tests online and just get help here on their tests even.

OpenStudy (anonymous):

such is the tragic state of our education system

OpenStudy (anonymous):

i need yur help. uber were did you go?

OpenStudy (anonymous):

Right here. What's up?

OpenStudy (anonymous):

He was helping.. you were the one who left.

OpenStudy (anonymous):

sorrry ! i thought he had read my other post cuss i update all in different post

OpenStudy (anonymous):

Simplify the expression (4x − 3)(x + 5).this one is next.(:

OpenStudy (anonymous):

Ok. Do you know about "FOIL"?

OpenStudy (anonymous):

yess, kinda. buht can you give me a brief summary bout iht.

OpenStudy (anonymous):

Sure. FOIL = "first outer inner last" This means you multiply each of the numbers in the brackets as given. So first, what is the first number in each parenthesis (including the x)?

OpenStudy (anonymous):

Very brief summary: \[(a+b)(c+d) = (a\times c) + (a\times d) + (b\times c) + (b \times d)\]

OpenStudy (anonymous):

Nice! I like that.

OpenStudy (anonymous):

I never do it that way myself.. I always just distribute.

OpenStudy (anonymous):

BTW, Polpak: I got bored enough to do 'normal' multiplication on polynomials, and found out it still works. I've always wondered why they make us do FOIL instead of, say, 3x + 7 * 2x - 4 ------- ^^^^^^ Unrelated, btw to your question jacyy, so ignore it :P ^^^^^^^

OpenStudy (anonymous):

ha ok. buht the thing i dont get ios if we are solving for ex or whatt?

OpenStudy (anonymous):

lol.. I always just do: (a + b)(c+d) = a(c+d) + b(c+d) Or (depending on what looks nicer) =c(a+b) + d(a+b)

OpenStudy (anonymous):

We aren't solving for x, just expanding out the expression.

OpenStudy (anonymous):

by carrying out the multiplication.

OpenStudy (anonymous):

Nope. We're just making a nicer looking equation. Getting rid of the parentheses basically.

OpenStudy (anonymous):

Like going from 4+4+4+4+4+4 to (4)*(6). Kinda.

OpenStudy (anonymous):

Alright, so first we need to make sure you understand FOIL. Do you know what it's trying to say?

OpenStudy (anonymous):

ohhh okaay, thanks. is the answer 4x^2+17x-5

OpenStudy (anonymous):

yess ii do understand foil:)

OpenStudy (anonymous):

Almost..

OpenStudy (anonymous):

Honestly, I don't think so. But let me double check. 4x − 3)(x + 5)

OpenStudy (anonymous):

Double check your math on the 'L'ast term.

OpenStudy (anonymous):

VERYYYYYYYY close, as polpak said. :D

OpenStudy (anonymous):

ARgghhh... is anyone else getting like 4 second lag between posts?

OpenStudy (anonymous):

okay will iht be - than + add the end

OpenStudy (anonymous):

no, look at your last terms: -3 and 5 What is -3*5 ?

myininaya (myininaya):

yeah i don't like teaching people about foil unless i thought them why you can do it (a+b)(c+d)=a(c+d)+b(c+d) or (a+b)c+(a+b)d

OpenStudy (anonymous):

yes, silly algebra teachers think that learning 'foil' is easier than distributing when you already went to the trouble of learning to distribute before foil.

myininaya (myininaya):

you can'd do foil on (a+b+c)(d+e+f) you have to distribute so there is almost no point in learning foil sense it is a weak method

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

(a+b)(c+d) foil only works on things like above

OpenStudy (anonymous):

Right. So 4x^2 + 17x -15

OpenStudy (anonymous):

Factor completely: 3x2 − 17x − 28 i got the other one thanks

OpenStudy (anonymous):

@mini: even then, I like doing classic multiplication as opposed to distributing. So much easier! :D I mean, I know it's still distributing, but it's so much more easier.

OpenStudy (anonymous):

So factoring is just the opposite.

OpenStudy (anonymous):

3x2 − 17x − 28 For this one, using the quadratic equation is the easiest way. Though the other folks can probably teach you the "real" way to do it.

OpenStudy (anonymous):

We are looking for some a and b such that: (3x + a)(x+b) = 3x^2 - 17x - 28

OpenStudy (anonymous):

Sadly, I am in Cal 2 and have yet to learn how to factor complex polynomials. Heh.

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

So using our 'foil' method: \[3x^2 + (3b + a)x + ab =3x^2 - 17x - 28\]

OpenStudy (anonymous):

Hmm... let's see if I can handle that...

OpenStudy (anonymous):

3x^2 + (3b+a)x + (4,7 or 14,2)

OpenStudy (anonymous):

plus or minus, of course.

myininaya (myininaya):

you can'd do foil on (a+b+c)(d+e+f) you have to distribute so there is almost no point in learning foil sense it is a weak method

myininaya (myininaya):

Ax^2+Bx+C find two factors of A*C that multiply to be A*C and add up to be B. A=3, B=-17, C=-28 A*C=3(-28)=12(-7)=6(-14)=2(-42)=4(-21) B=-17=-21+4 so replace middle term with Bx=-17x with -21x+4x so we have 3x^2-17x-28 3x^2-21x+4x-28 then factor by grouping 3x(x-7)+4(x-7) (x-7)(3x+4)

OpenStudy (anonymous):

Oh snap. Is it 3x^2 + x(3*7-4) + 7*-4?

OpenStudy (anonymous):

Oh sorry, missed your post. And also realized I got positive 17. Bleh.

OpenStudy (anonymous):

sorry i just came back

OpenStudy (anonymous):

Question: 13 of 40: Factor completely: 3x2 − 17x − 28Select one of the options below as your answer: A. (3x + 4)(x + 7) B. (3x + 4)(x − 7) C. 3x2 − 17x − 28 D. (3x2 + 4)(x − 7) thats the question and answer s

OpenStudy (anonymous):

That means: 3b+a = -17 and a*b = -28

OpenStudy (anonymous):

and I see the others have already done it.

myininaya (myininaya):

you know multiplication is communative right? a*b=b*a 3*2=2*3

myininaya (myininaya):

commutative*

OpenStudy (anonymous):

im confused:/

myininaya (myininaya):

which part confuses you?